An automobile accelerates uniformly from rest and reaches a velocity of 22 m/s in 9 s. The tire  diameter is 58 cm. Determine the number of revolutions the tire makes during this motion  and the final angular velocity of the tire in revolutions per second.



1
Expert's answer
2021-03-10T17:21:36-0500

a) Let's first find the acceleration of the automobile:


"a=\dfrac{v-v_0}{t}=\dfrac{22\ \dfrac{m}{s}}{9\ s}=2.44\ \dfrac{m}{s^2}."

Then, let's find the distance covered by the automobile during this motion:


"s=\dfrac{1}{2}at^2=\dfrac{1}{2}\cdot2.44\ \dfrac{m}{s^2}\cdot(9\ s)^2=99\ m."

Let's find the distance covered by the tire in one revolution:


"d=C=\pi d=3.14\cdot0.58\ m=1.82\ m."

Finally, we can find the number of revolutions the tire makes during this motion:


"n=\dfrac{s}{C}=\dfrac{99\ m}{1.82\ \dfrac{m}{rev}}=54.4\ rev."

b) We can find the final angular velocity of the tire as follows:


"v=\omega r,""\omega=\dfrac{v}{r}=\dfrac{22\ \dfrac{m}{s}}{0.29\ m}=75.9\ \dfrac{rad}{s},""\omega=75.9\ \dfrac{rad}{s}\cdot\dfrac{1\ rev}{2\pi\ rad}=12.1\ \dfrac{rev}{s}."